4. Thirteen telephones have just been received at a cell phone repair service and are randomly assigned an order to be repaired. Four of these are iPhones, Seven are Samsung phones, and the other two are LG phones including one that is red. a) What is the probability that two cellphones of each type (iPhone, Samsung, and LG) are among the first six phones serviced? b) What is the chance that the next phone serviced after the first LG phone is the red LG phone? (Hint. see example 5j in section 2.5 of the text) c) What is the probability that at least one LG is amongst the first six phones given that four of these six are Samsung phones.
Answer #1

a)

We have to find the Probability that two cell phones of each type are among first six phone serviced

Total number of ways of drawing six phones out of 15 are 13C6

131 13 × 12 × 11 × 10 × 9 × 8×7! 617! 13 × 12 × 11 × 10×9×8×7! 720 x 7 13C6 617!

1716

We have to select 6 different phones & two cell phones of each type (iphone,samsung and LG)

Total Number of ways for doing this

4! (4C2)X (7C2)X22)X 212 25 2

41-71-21 212 25 2 4×3×21.7×6××5! (4C2)X(7C2)X(2C2) =-X 215!

6 × 21 × 1 126 ways

Probability will be given as below:

1260.0734 1716

b)

Let E1 be the event that first phone serviced was LG.

Let E2 be the event that after first LG phone serviced second is also LG.

We have to find- probability of event E2, after occurrence of event E1.

P(E1) will be

Selecting one phone among 2 phones from total of 13 phones.

P(E1) = (2C1) / 13C1

P(E1) = 2/13

Because the phone will not be replaced

hence the number of phones left to be serviced are 12.

Hence, the probability the second phone is LG is-

P(E2) = (1C1) / 12C1

P(E2) = 1/12

The probability of next phone serviced after the first LG phone is the red LG phone-

P(E)= P(E1)*P(E2)=(2/13)x(1/12)=2/156=0.0128

c)

This can be solved by taking conditional probabilities.

Let E1 be the event that at least one phone is LG among six selected.

Let E2 be the event that four of six are Samsung .

So here our event of interest is P(E1|E2).

And P(E1|E2) = P(E1 and E2)/P(E2)

In this case, we can say that both the events are independent, like at least one phone among six is LG will nothing to do with four phones among six are Samsung.

So, in case of independent events- conditional probability will be-

P(E1|E2) = P(E1)

So, probability of at least one LG among six phones selected.

= 1- P(None is LG) = 1- 11C6 / 13C6=1-462/1716=1-0.269=0.73 ans

(Because there are 4 iphones and 7 samsung phones =11 which are not LG)

Probability of at least one LG will be 0.73